Suppose C=50+0.80Y, Y=Disposable income,Investment(I)=70,Government expenditure(G)=70,Tax(Tx)=0.10Y.
a.Find equilibrium output.
b.What are tax revenues at equilibrium output?Does the government have a balanced budget?
c.Find equilibrium output when investment increases from 70 to 90.
1
Expert's answer
2015-01-14T03:20:13-0500
We can write an equation
a.Y = Equilibrium output = C + I + G
Y = 50+0.80Y + 70 + 70
Y = 190 + 0.80Y
Y = 950

b. Tax(Tx)=0.10Y = 0.10 * 950 = 95
Tax(Tx) > Government expenditure(G), so the governmenthave budget proficit

c. Y’ = Y + (90 -70) = 950 + 20 = 970